OpenStudy (anonymous):

find the angle between the given vectors to the nearest tenth of a degree u=<-5,8> v=<-4,8>? can someone please help...

OpenStudy (anonymous):

\[u=-5i+8j,v=-4i+8j,u.v=\left( -5i+8j \right).\left( -4i+8j \right)=\left( -5 \right)\left( -4 \right)+\left( 8 \right)\left( 8 \right)\] =20+64=84 Let theta be the angle between two vectors.Then \[u.v=\left| u \right|\left| v \right|\cos \theta \] can you complete it ?

OpenStudy (anonymous):

Well that's the only part I understand how to do... the rest is what I don't understand @surjithayer